OpenStudy (anonymous):

Solve the equation for y. 8x – 9y = 11

OpenStudy (anonymous):

8x-9y=11 9y=8x-11 y=8/9x-11/9

OpenStudy (anonymous):

that's a choice

OpenStudy (anonymous):

????

OpenStudy (anonymous):

help again? @ayoubEpst

OpenStudy (anonymous):

I will make new quiestion

OpenStudy (anonymous):

i give you the full answer